# Interplay between strain and size quantization in a class of topological insulators based on inverted-band semiconductors

### v1

@@ V Gap in the V-P states for the finite-size sample
Above we have considered several problems with the use of the boundary conditions Eq. (21) in the case of a single interface. We have found, in particular, the location of the Dirac point of the Volkov-Pankratov states as a function of the bands’ parameters of the neighbouring materials. A gap is expected to open in the VP states in the case when a film width is finite because of a hybridization of the states belonging to the opposite boundaries. This quantity is also important for the interpretation of the ARPES data. We now consider the effects of finite size on the gap value and spectrum of the VP states. We will be interested in the VP states spectra for small values of $k_{y}$, at least $k_{y}\ll E_{g}/P$, thus we can neglect the interaction with the heavy holes. For simplicity we consider also the zero strain case. Under these conditions, from Eq. (2) of Ref. Khaetskii, we obtain the following two-band model Hamiltonian
@@
For the film of the width $d$ with the high barriers one obtains the following equation for the V-P surface states spectrum
@@
It is interesting that the velocities of the V-P branches depend on the parameter $\eta$ which describes the asymmetry of the offsets in the conduction and valence bands. In the limit of large width $(\alpha/s\gg 1)$, the solution is $\xi\to\xi_{\infty}$ and the spectrum is
@@
Thus in this limit the gap value is exponentially small with respect to the film width. For the derivation of Eq. (41) see Appendix A. Eq. (41) shows that reducing $d$ leads to the opening of a gap in these states with the upper branch increasing in energy and the lower branch decreasing, see also Fig. 3 of the Ref. Subashiev, . One can say that the location of the central point of the V-P branches stays at rest within the first-order approximation, but a gap opens between the two Dirac cones, see Appendix A.
It is important to realize what happens near the critical width $d_{1}$ when the upper branch of the Volkov-Pankratov states ($S1$ branch in Fig. 4 of Ref. Subashiev, ) is inverted with the $hh1$ sub-band, which is the first quantized sub-band of the heavy holes. Near this width the energy spectrum (Fig. 4(b) of Ref. Subashiev, ) will look like a linear-in-$k_{y}$ Dirac dependence with no gap (see also Ref. Dietl, ). One might erroneously conclude that near this energy one has Volkov-Pankratov states with DP close to the $\Gamma_{8}$ point. To the contrary, the lower branch of the Volkov-Pankratov states (i.e. $S2$ branch, see Figs. (3,4) of Ref. Subashiev, ) lies well below this energy, and the DP also lies below. Note that at $d>d_{1}$ (see Fig. 4(b) of Ref. Subashiev, ) the inverted $hh1$ state coincides with the DK1 surface state at $k_{y}d>\pi$ For the details see Ref. Dyak1, .
@@
VI Interplay between the finite size effects and applied strain (surface states within Luttinger model)
We consider here some physics related to the interplay between the finite size effects and applied strain. The important questions can be answered already within the Luttinger model, while considering the characteristic energies near the $\Gamma_{8}$ point. Moreover, the essential physics can be studied within the model with high barriers and zero boundary conditions (i.e. the exact form of the BCs is not important). Everywhere in this section the origin of the energy is at the $\Gamma_{8}$ point. We consider a film of the width $d$ with a center of inversion (identical boundaries). We note that all solutions are doubly degenerate with respect to spin because of inversion symmetry. As a measure of topological protection of the surface states at opposite boundaries, one can choose to consider $1-J$, where $J$ is the following overlap integral
@@ VI.1 Positive (biaxially tensile in-plane) strain, $\Delta>0$
As a result, the DK1 branch is transformed into the first heavy-hole subband $hh1$, see Fig. 3. This transformation occurs at the $k_{y}=k_{y}^{\star}$ value which is found from the condition $\kappa_{l}=0$ (the expression for $\kappa_{l}$ and the exact equation for $k_{y}^{\star}$ obtained from this condition are given in Appendix B by Eq. (49) and Eq. (48), respectively). At this point the DK1 state crosses the edge of the bulk heavy-hole band. One can also say that at the point $k_{y}^{\star}$ the heavy-hole components of the $hh1$ state transform from the form $\sin k_{x}x,\cos k_{x}x$ (localized near the center of the film) to the exponents which are localized more near the boundaries (i.e. $hh1$ bulk-like state transforms into the DK1 surface state), see also Ref. Dyak1, .
Allowed values of transverse momenta at $k_{y}=0$ are $k_{x}^{(n)}=\pi n/d$ and the corresponding energies of the light and heavy-hole subbands are $E_{l}^{(n)}=\hbar^{2}\pi^{2}n^{2}/2m_{l}d^{2}+\Delta/2$ and $E_{h}^{(n)}=-\hbar^{2}\pi^{2}n^{2}/2m_{h}d^{2}-\Delta/2$, respectively, see Fig. 3.
The parameter which determines the crossover between the regimes with an exponentially small value of $J$ and the value of $J\simeq 1$ is $\theta_{d}=d\sqrt{m_{l}\Delta}/\pi\hbar$. (Everywhere in this section we assume $\beta=m_{l}/m_{h}\ll 1$). At large width $d$ such that $\theta_{d}\gg 1$ Fig. 3 (a) describes a material which still can be treated as a 3D topological insulator. For that, the Fermi level should lie within the $hh1$ branch above the edge of the bulk heavy-hole band (shown by the dashed line) and below the bottom of the $lh1$ sub-band. There are no bulk states at the Fermi level for this situation, besides the overlap $J$ of the surface DK1 states belonging to the opposite boundaries is exponentially small since $\theta_{d}\gg 1$. This case of a big strain corresponds to the red curve of Fig. 4, where the overlap integral $J$ of two DK1 surface states localized near the opposite surfaces and having opposite $k_{y}$ vectors is plotted as a function of $k_{y}$ for HgTe. The maximally allowed value of $k_{y}$ ($k_{y}^{\text{max}}$) corresponding to the energy of the bottom of the $lh1$ sub-band is shown by the red arrow. We see that for parameters corresponding to the red curve (which can be realized for realistic values of strain only for relatively wide films $d>70nm$) one has the lowest value of the overlap integral $J\approx 0.02$.
[…]
@@
Note that with increasing the positive strain value the overlap integral increases (the red curve versus a blue one) which is quite a counterintuitive behavior. To understand it we note that in the case of exponentially small overlap one has the following analytical expression for $J$ (in the case of $\beta\ll 1$)
@@
We have obtained the above expression for the decaying factor of light holes $\kappa_{l}$ using Eqs. (26, 27) of Ref. Khaetskii, . Thus we see that the decaying vector of light holes decreases with strain, in opposite to what might naively expect.
@@
The general conclusion in the case of positive strain is that for the film widths within the usual experimental range of $10nm<d<100nm$ and for realistic strain values the overlap integral of the DK1 states belonging to the opposite boundaries is not small and spin conserving backscattering between them is not suppressed. As a result, it will destroy the robustness of the surface state.
[…]
@@ VI.2 Negative (biaxially compressive in-plane) strain, $\Delta<0$
Let us consider biaxial, compressive in-plane strain (in accord with the definition we used in Ref. Khaetskii, it corresponds to the strain-induced energy $\Delta<0$). Then for large enough sample without size-quantization effects we have a Dirac semi-metal (DSM) phase. We consider here the transition between the DSM phase and the quasi 3D (quantized) topological insulator phase (as it is called in the literature Coster ; Vail ) through the reduction of the layer width $d$. We will argue that below some critical width $d_{|\Delta|}$ the DSM transforms into an insulator which, strictly speaking, cannot be considered as topological for reasons similar to the ones described in the previous subsection. We consider the width in the interval
@@
where $d_{1}$ is the width at which the upper branch of the Volkov-Pankratov states becomes inverted with the first quantized sub-band of the heavy holes $hh1$, see Ref. Subashiev, . $d_{|\Delta|}$ is the characteristic sample width at which the strain-induced energy is equal to the sum of the quantization energies of the light and heavy holes. (In the case of HgTe and for a realistic strain value corresponding to $\Delta=20meV$ the length $d_{|\Delta|}\approx 26nm$). The left inequality in Eq. (44) means that the width is already large enough so that both branches of the Volkov-Pankratov states lie well below the energy of the $\Gamma_{8}$ point and we can describe the problem within the Luttinger model. The right inequality in Eq. (44) means that the width is small enough to “quench” the DSM state (i.e. the energy due to the size quantization is larger than the original overlap of the light and heavy-hole bands due to strain), and the transition to a new state has already happened. Our task is to understand the properties of this state. The physical pattern presented in recent publications (see, for example, Ref. Vail, ) is unclear, especially the nature of ”surface” and ”bulk” gaps discussed there.
@@
We argue that the physics for this situation was described in Refs. (Dyak1, ,Subashiev, ). Though in both papers strain was assumed to be zero, this fact does not make a qualitative difference as far as we consider a situation when an original overlap of the light and heavy-hole bands is overcome by the size quantization. The only difference is that due to a ”compensation” of the negative strain by the quantized energies the energy difference between the bottoms of the $lh1$ and $hh1$ sub-bands can be small if one considers the situation close to the transition point (where $d=d_{|\Delta|}$). Such a situation is assumed in Fig. 5, where several lowest sub-bands (closest to zero energy) are presented. Once again, due to the inversion, the $hh1$ branch increases with energy instead of decreasing and at large enough $k_{y}$ transforms into the DK1 surface state.
@@
We claim that Fig. 5 corresponds to the situation presented in Fig. 2(c) in Ref. Vail, . (We note that Fig. 2(b) of Ref. Vail, demonstrates indeed that the width considered is already large enough, so that both branches of the Volkov-Pankratov states lie well below). The authors of Ref. Vail, named the distance between the bottoms of the $lh1$ and $hh3$ sub-bands as ”the bulk state band gap” and the distance between the bottoms of the $hh1$ (DK1) and $hh2$ sub-bands as ”the surface state band gap”. We see that these quantities actually have different physical meaning. We would say that the former quantity does not have any special physical meaning, but the latter should rather be named as ”the bulk gap” of this quasi 3D system, see also Ref. Dyak1, .
@@
Assume now that the conditions of Eq. (44) are fulfilled, and, in addition, the Fermi level is within the $hh1$ (DK1) state, see Fig. 5. Then despite the fact that the Fermi level lies below the bottom of the first light-hole sub-band ($lh1$), this DK1 state cannot be treated as a truly topological one since there is an appreciable overlap of the surface states wave functions located mostly at opposite boundaries (with the same spin orientations and opposite $k_{y}$). The reason for that is the fact that while keeping the Fermi level location below the bottom of the $lh1$ sub-band one cannot fulfill the condition $k_{y}d\gg\pi$, see the discussion in the previous subsection. As a result, an elastic impurity scattering which keeps the spin orientation but changes the sign of $k_{y}$ will destroy the robustness of the surface state.
@@
VII Conclusions

### v2

@@ V Gap in the Volkov-Pankratov states
Above we have found the location of the Dirac point of the Volkov-Pankratov states as a function of the bands’ parameters of the neighbouring materials in the case of a single interface. A gap is expected to open in the VP states in the case when a film width is finite because of a hybridization of the states belonging to the opposite boundaries. This quantity is also important for the interpretation of the ARPES data. We now consider the effects of a sample finite size on the gap value and spectrum of the VP states. We will be interested in the VP states spectra for small values of $k_{y}$, at least $k_{y}\ll E_{g}/P$, thus we can neglect the interaction with the heavy holes. For simplicity we consider also the zero strain case. Under these conditions, from Eq. (2) of Ref. Khaetskii, we obtain the following two-band model Hamiltonian
@@
For the film of the width $d$ with the high barriers using the boundary conditions Eq. (39) one obtains the following equation for the VP surface states spectrum
@@
It is interesting that the velocities of the VP branches depend on the parameter $\eta$ which describes the asymmetry of the offsets in the conduction and valence bands. In the limit of large width $(\alpha/s\gg 1)$, the solution is $\xi\to\xi_{\infty}$ and the spectrum is
@@
Thus in this limit the gap value is exponentially small with respect to the film width. For the derivation of Eq. (21) see Appendix C. Eq. (21) shows that reducing $d$ leads to the opening of a gap in these states with the upper branch increasing in energy and the lower branch decreasing, see also Fig. 3 of the Ref. Subashiev, . One can say that the location of the central point of the VP branches stays at rest within the first-order approximation, but a gap opens between the two Dirac cones, see Appendix C.
It is important to realize what happens near the critical width $d_{1}$ when the upper branch of the Volkov-Pankratov states ($S1$ branch in Fig. 4 of Ref. Subashiev, ) is inverted with the $hh1$ sub-band, which is the first quantized sub-band of the heavy holes (see also Ref. Bernevig1, ). Near this width the energy spectrum (Fig. 4(b) of Ref. Subashiev, ) will look like a linear-in-$k_{y}$ Dirac dependence with no gap (see also Ref. Dietl, ). One might erroneously conclude that near this energy one has Volkov-Pankratov states with DP close to the $\Gamma_{8}$ point. To the contrary, the lower branch of the Volkov-Pankratov states (i.e. $S2$ branch, see Figs. (3,4) of Ref. Subashiev, ) lies well below this energy. Therefore, the ”DP” (i.e. the energy where both $S1$ and $S2$ branches match each other in the limit of a very large thickness) also lies below. Note that at $d>d_{1}$ (see Fig. 4(b) of Ref. Subashiev, ) the inverted $hh1$ state coincides with the DK1 surface state at $k_{y}d>\pi$ For the details see Ref. Dyak1, .
@@
VI Interplay between the finite size effects and applied strain
As we have already mentioned in the Introduction, the most important thing in the description of the transitions related to the competition between the finite size effects and applied strain is the correct identification of the subbands involved in the corresponding transitions. It is not enough just to find the width of the film where some subbands cross each other and claim that it is the point of the transition. One needs also to investigate the dispersions of the corresponding subbands as the function of the in-plane momentum. Depending on the character of these dispersions, whether they are of the normal or an inverted type, the description of the physics is very different. For example, the authors of Refs. (Coster, ; Anh, ; Vail, ) did not take this fact into account which led them, in particular, to an incorrect conclusion about the physical meaning of various gaps formed in the system under the size quantization, see below.
Moreover, none of the papers Refs. (Ohtsubo, ; Coster, ; Anh, ; Vail, ) investigates a robustness of the topological surface states near the transition point, which is determined by the overlap of the wave functions of the surface states located near the opposite boundaries. This issue is considered for the first time in the current work. We have shown (see discusson below and Fig. 4) that the parameters of the problem are such that an appreciable overlap of the surface states’ wave functions located at opposite boundaries occurs. As a result, an elastic (and spin-conserving) impurity scattering between the states located at opposite boundaries will destroy the robustness of the surface state.
In the following, we consider a film of the width $d$ with a center of inversion (identical boundaries). Subsections A and B below work through the positive and negative cases of strain, respectively.
The important questions regarding the transitions considered in this section can be answered already within the Luttinger model, while considering the characteristic energies near the $\Gamma_{8}$ point. Moreover, the essential physics can be studied within the model with high barriers, then the BCs reduce to the zero values for both light-hole and heavy-hole components; see the corresponding discussion in Appendix A, after Eq. (40). Everywhere in this section the origin of the energy is at the $\Gamma_{8}$ point. We note that all solutions are doubly degenerate with respect to spin because of inversion symmetry. As a measure of topological protection of the surface states at opposite boundaries, one can choose to consider $1-J$, where $J$ is the following overlap integral
@@ VI.1 Positive (biaxially tensile in-plane) strain, $\Delta>0$
As a result, the DK1 branch is transformed into the first heavy-hole subband $hh1$, see Fig. 3. This transformation occurs at the $k_{y}=k_{y}^{\star}$ value which is found from the condition $\kappa_{l}=0$ (the expression for $\kappa_{l}$ and the exact equation for $k_{y}^{\star}$ obtained from this condition are given in Appendix D by Eq. (53) and Eq. (52), respectively). At this point the DK1 state crosses the edge of the bulk heavy-hole band. One can also say that at the point $k_{y}^{\star}$ the heavy-hole components of the $hh1$ state transform from the form $\sin k_{x}x,\cos k_{x}x$ (localized near the center of the film) to the exponents which are localized more near the boundaries (i.e. $hh1$ bulk-like state transforms into the DK1 surface state), see also Ref. Dyak1, .
Allowed values of transverse momenta at $k_{y}=0$ are $k_{x}^{(n)}=\pi n/d$ and the corresponding energies of the light and heavy-hole subbands are $E_{\textsl{lhn}}=\hbar^{2}\pi^{2}n^{2}/2m_{l}d^{2}+\Delta/2$ and $E_{\textsl{hhn}}=-\hbar^{2}\pi^{2}n^{2}/2m_{h}d^{2}-\Delta/2$, respectively, see Fig. 3.
The parameter which determines the crossover between the regimes with an exponentially small value of $J$ and the value of $J\simeq 1$ is $\theta_{d}=d\sqrt{m_{l}\Delta}/\pi\hbar$. (Everywhere in this section we assume $\beta=m_{l}/m_{h}\ll 1$). At large width $d$ such that $\theta_{d}\gg 1$, Fig. 3 (a) describes a material which still can be treated as a 3D topological insulator. For that, the Fermi level should lie within the $hh1$ branch above the edge of the bulk heavy-hole band (shown by the dashed line) and below the bottom of the $lh1$ sub-band. There are no bulk states at the Fermi level for this situation, furthermore the overlap $J$ of the surface DK1 states belonging to the opposite boundaries is exponentially small since $\theta_{d}\gg 1$. This case of a big strain corresponds to the red curve of Fig. 4, where the overlap integral $J$ of two DK1 surface states localized near the opposite surfaces and having opposite $k_{y}$ vectors is plotted as a function of $k_{y}$ for HgTe. The maximally allowed value of $k_{y}$ ($k_{y}^{\text{max}}$) corresponding to the energy of the bottom of the $lh1$ sub-band is shown by the red arrow. We see that for parameters corresponding to the red curve (which can be realized for realistic values of strain only for relatively wide films $d>70nm$) one has the lowest value of the overlap integral $J\approx 0.02$.
For thinner films the value of $J$ will be much larger. The red curve will become closer to the blue one corresponding to zero strain, and $k_{y}^{\text{max}}$ from being proportional to $\sqrt{\Delta}$ at large $\Delta$ (and large $d$) will change with decreasing $d$ tending to $\simeq\pi/d$. For example, for a width $40nm$ and the strain value corresponding to $\Delta=20meV$ the parameter $\Delta/\delta_{h}$ is around 40, and the overlap integral value for a maximally allowed $k_{y}^{\text{max}}$ is $J\approx 0.25$. (Here $\delta_{\textit{h}}$ is the heavy-hole quantization energy, $\delta_{\textit{h}}=\hbar^{2}\pi^{2}/2m_{h}d^{2}$).
@@
Observe that with increasing the positive strain value the overlap integral increases (the red curve versus a blue one) which is quite a counterintuitive behavior. To understand it we note that in the case of exponentially small overlap one has the following analytical expression for $J$ (in the case of $\beta\ll 1$)
@@
We have obtained the above expression for the decaying factor of light holes $\kappa_{l}$ using Eqs. (26, 27) of Ref. Khaetskii, . Thus we see that the decaying vector of light holes decreases with strain, in opposition to what might naively expect.
@@
The general conclusion in the case of positive strain is that for the film widths within the usual experimental range of $10nm<d<100nm$ and for realistic strain values the overlap integral of the DK1 states belonging to the opposite boundaries is not small and spin-conserving backscattering between them is not suppressed. As a result, it will destroy the robustness of the surface state.
Figure 4: The overlap integral $J$ in Eq. (22) calculated for the $hh1$ (DK1) subband in a HgTe quantum well for different values of strain. The strain parameter $\Delta$ is measured in units of the heavy-hole quantization energy $\delta_{\textit{h}}=\hbar^{2}\pi^{2}/2m_{h}d^{2}$. (a) The filled circles show the overlap obtained from a numerical diagonalization of the Luttinger Hamiltonian, whereas the solid lines show the result of the weak-overlap approximation, which is continued by dashed lines in the regime where the approximation breaks down. The arrows with dash-dotted vertical lines in (a) show the values $k_{y}^{\text{max}}$ at which the bottom of the bulk-like $lh1$-subband equals to the energy of the surface-like DK1 state. (b) Same as in (a), but on a logarithmic scale. Parameters of HgTe: $m_{h}=0.48\,m_{0}$ and $m_{l}=0.03\,m_{0}$.
@@ VI.2 Negative (biaxially compressive in-plane) strain, $\Delta<0$
Let us consider biaxial, compressive in-plane strain. In accord with the definition we used in Ref. Khaetskii, it corresponds to the strain-induced energy $\Delta<0$. Then for a large enough sample without size-quantization effects we have a Dirac semi-metal (DSM) phase. We consider here the transition between the DSM phase and the quasi 3D (quantized) topological insulator phase (as it is called in the literature Coster ; Vail ) through the reduction of the layer width $d$. We will argue that below some critical width $d_{|\Delta|}$ the DSM transforms into an insulator which, strictly speaking, cannot be considered as topological for reasons similar to the ones described in the previous subsection. We consider the width in the interval
@@
where $d_{1}$ is the width at which the upper branch of the Volkov-Pankratov states becomes inverted with the first quantized sub-band of the heavy holes $hh1$, see Ref. Subashiev, . $d_{|\Delta|}$ is the characteristic sample width at which the strain-induced energy is equal to the sum of the quantization energies of the light and heavy holes. (In the case of HgTe and for a realistic strain value corresponding to $\Delta=20meV$ the length $d_{|\Delta|}\approx 26nm$.) The left inequality in Eq. (24) means that the width is already large enough so that both branches of the Volkov-Pankratov states lie well below the energy of the $\Gamma_{8}$ point and we can describe the problem within the Luttinger model. The right inequality in Eq. (24) means that the width is small enough to “quench” the DSM state (i.e. the energy due to the size quantization is larger than the original overlap of the light and heavy-hole bands due to strain), and the transition to a new state has already happened. Our task in this section is to understand the properties of this state. The physical model presented in recent publications (see, for example, Ref. Vail, ) is unclear, especially the nature of ”surface” and ”bulk” gaps discussed there.
@@
We argue that the physics for this situation was described in Refs. (Dyak1, ,Subashiev, ). Though in both papers strain was assumed to be zero, this fact does not make a qualitative difference when an original overlap of the light and heavy-hole bands is overcome by the size quantization. The only difference is that the energy due to the quantization compensates the negative strain energy such that the difference between the minima of the $lh1$ and $hh1$ sub-bands is small close to the transition point where $d=d_{|\Delta|}$. Such a situation is assumed in Fig. 5, where several lowest sub-bands (closest to zero energy) are presented. Once again, due to the inversion, the $hh1$ branch increases with energy instead of decreasing and at large enough $k_{y}$ transforms into the DK1 surface state.
@@
We claim that Fig. 5 corresponds to the situation presented in Fig. 2(c) in Ref. Vail, . (We note that Fig. 2(b) of Ref. Vail, demonstrates indeed that the width considered is already large enough, so that both branches of the Volkov-Pankratov states lie well below $\Gamma_{8}$ point). The authors of Ref. Vail, named the distance between the bottoms of the $lh1$ and $hh3$ sub-bands as ”the bulk state band gap” and the distance between the bottoms of the $hh1$ (DK1) and $hh2$ sub-bands as ”the surface state band gap”. We find that these quantities actually have different physical meaning. We would say that the former quantity does not have any special physical meaning, but the latter should rather be named as ”the bulk gap” of this quasi 3D system, see also Ref. Dyak1, .
@@
Assume now that the conditions of Eq. (24) are fulfilled, and, in addition, the Fermi level is within the $hh1$ (DK1) state, see Fig. 5. Then despite the fact that the Fermi level lies below the bottom of the first light-hole sub-band ($lh1$), this DK1 state cannot be treated as a truly topological one since there is an appreciable overlap of the surface states wave functions located mostly at opposite boundaries (with the same spin orientations and opposite $k_{y}$). The reason for that is the fact that while keeping the Fermi level location below the bottom of the $lh1$ sub-band one cannot fulfill the condition $k_{y}d\gg\pi$ (see the discussion in the previous subsection). As a result, an elastic impurity scattering which keeps the spin orientation but changes the sign of $k_{y}$ will destroy the robustness of the surface state.
@@
VII Discussions
Next we discuss our results in the context of the work in Refs. Ohtsubo, ; Coster, ; Anh, ; Vail, , where different transitions with varying $d$ were studied both experimentally and theoretically.
[…]
One could envision a scenario in which a large built-in electric field forms in the ultrathin $\alpha$-Sn film due to inequivalence of surfaces and associated charge transfers. In that case, the description of the splitting of the VP states derived in Sec. V is modified. In essence, if the bias due to the electric field exceeds the splitting $\Delta_{\pm}$ given in Eq. (21), then two Dirac cones form, one per surface, with Dirac points shifted with respect to each other by the value of the bias. The splitting between the VP states on opposite surfaces occurs then at a finite in-plane momentum. The overall spectrum resembles two Rashba spectra, one dispersing upwards in energy and one downwards. Each of these Rashba-like spectra has a Dirac cone at zero in-plane momentum. If the overlap integral is sufficiently small, the states detected on one of the two surfaces may appear as a pristine Dirac cone, because the states on the opposite surface are not visible in ARPES. Such an off-resonance scenario of interaction between the VP states on opposite surfaces might explain the unusual robustness of the Dirac cone observed in the experiment of Ref. Ohtsubo, .
In Ref. Anh, , epitaxial growth of $\alpha$-Sn on InSb with high quality is presented. The studied samples exhibit unprecedentedly high quantum mobilities of the surface states (30 000 $\text{cm}^{2}/\text{V}s$), which is ten times higher than the previously reported values. The authors study the interfacial and bulk band structure of $\alpha$-Sn via the Shubnikov–de Haas oscillations combined with the first-principles calculations and claim that the results identify that $\alpha$-Sn grown on InSb is a topological Dirac semimetal (TDS). Furthermore, a crossover from the TDS to a 2D topological insulator and a subsequent phase transition to a trivial insulator when varying the thickness of $\alpha$-Sn are demonstrated.
We remark that the DSM–quantized 3D TI transition, which is one of the main interests of the current work, is not mentioned in Ref. Anh, . From the ab initio modelling of the $\alpha$-Sn/InSb structure, a very strong compressive in-plane strain for the $\alpha$-Sn layer is obtained in Ref. Anh, . The strain value is five times larger than the value used by other research teams dealing with this system. Nonetheless, the strain does not seem to affect significantly the sub-band energy separation, or at least in ways that are expected from the Kane model. Thus, in Fig. S8 of the Supporting Information of Ref. Anh, , the sub-band structure for the 56 ML-thick $\alpha$-Sn films under different in-plane strain (from strongly tensile to strongly compressive ones) is presented. The energy separation between the bottoms of the two lowest sub-bands with positive energies (which in our classification are $hh1$ and $lh1$) stays practically constant as a function of strain. From the Kane model, we expect that energy separation to decrease to half of its value in going from Fig. S8 (a) to Fig. S8 (c). Recall that the gross effect of the strain is to shift the light-hole spectrum with respect to the heavy-hole spectrum, since the two sectors separate from each other at zero in-plane momentum. It is relatively easy to distinguish between a sparse ladder of light-hole sub-bands and a dense ladder of heavy hole sub-bands and assess this gross effect of the strain in produced sub-band structure. The strain effect obtained in Ref. Anh, seem to contradict this basic expectation.
[…]
In summary, our results and conclusions differ significantly from those presented in Refs. Ohtsubo, ; Coster, ; Anh, ; Vail, . These differences concern the correct interpretation of the sub-band structure in the quantum well as well as the identifications of all regimes in the presence of strain and quantization, and the transitions between them. In particular, we assessed the robustness of the topological surface states in the region of the crossover between the DSM and quantized 3D TI phases, see Fig. 4, which was not done before.
VIII Conclusions